OpenStudy (anonymous):

What is the third–degree polynomial function such that f(0) = –18 and whose zeros are 1, 2, and 3?

OpenStudy (anonymous):

Ahh thank you!!!

hero (hero):

wait, lol. I didn't see the f(0) = -18 part

OpenStudy (anonymous):

oh. so what do i do im soooo confused!

OpenStudy (anonymous):

you know it is \[f(x)=a(x-1)(x-2)(x-3)\] and you also know the constant is 18 since \[f(0)=18\] the only thing you don't know is "a"

OpenStudy (anonymous):

\[a\times -1\times -2\times -3=-6a=18\] therefore \[a=-3\]

OpenStudy (anonymous):

and so your answer is \[f(x)=-3(x-1)(x-2)(x-3)=-3 x^3+18 x^2-33 x+18\]

OpenStudy (anonymous):

@luis unfortunately when you tack on -18 at the end of the product, the zeros are no longer 1,2 and 3

hero (hero):

satellite is the master

OpenStudy (anonymous):

of disaster!

OpenStudy (anonymous):

lol i hate algebra!!!! its mean